If D(p)=100/p and c(y)=y2


what is the optimal level of output of the monopolist?


1
Expert's answer
2021-04-20T07:28:30-0400

In order to find the optimal level of output of the monopolist, marginal revenue (MR) must be equal to marginal cost (MC). That is, "MR=MC"

In this equation, we will use the inverse elasticity rule for the demand curve "D(p)=\frac{100}{p}". The demand curve has a constant elasticity of "-1".

Here is how we will calculate the marginal revenue.

"MR=q=\frac{100}{p}"


"p=\frac{100}{q}"


"Revenue(R)=p(q)q"


"R=\frac{100}{q} \times q=100"


"MR=\frac{dR}{dq}=0"


"MR=0"


Therefore, for all levels of output, the marginal revenue will be 0. "MR\neq MC"

This indicates that profits are decreasing. The company can then generate the smallest amount of production possible, as profits will decline as a result, and profits will be maximized.